# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v2

@@ 4 Discussions
  • Optimality. Based on the $\Omega(\sqrt{T})$ lower bound on the expected regret provided by Scarlett [43], we conjecture that our $\widetilde{O}(\sqrt{T})$ high-probability regret bound for GP-UCB is near-optimal. However, it is not straightforward to extend the lower bound for the expected regret in [43] to a high probability result. Specifically, the lower bound in [43] is quantified by a mutual information term (Lemma 4 in [43]); however, to our knowledge, the technique used to handle this term appears to be specific to the expected regret setting. We believe that the rigorous optimality argument for the Bayesian high probability regret is an important direction for future research.
@@
  • Smoothness condition. In our result for the Matérn kernel, we require an additional smoothness constraint to obtain a $\widetilde{O}(\sqrt{T})$ regret bound. To overcome this issue in our proof, we believe that we need stronger regularity conditions on the sample path around the maximizer than those assumed in Lemma 2.
@@
  • Extension to the expected regret. Our regret bounds involve regularity constants that depend on the sample path. However, to our knowledge, there is no existing research that rigorously analyzes how these constants depend on the confidence level $\delta_{\mathrm{GP}}$. This makes it difficult to obtain the expected regret guarantees as with the original GP-UCB, whose expected regret bounds are established by properly decreasing the confidence level as a function of $T$ (e.g., [37, 49]). To overcome this issue, further analysis for Lemma 2, or another idea to quantify the sample path regularities, is required.
@@
  • Extension to other algorithms. One limitation of our technique is its restricted applicability to other algorithms. To apply our proof, at least the algorithm should satisfy the following two conditions: (i) on any index subset, the sub-linear cumulative regret is obtained with high probability (Lemma 21), and (ii) the high probability lenient regret bound is provided (Lemma 6). Other major algorithms in the Bayesian setting (e.g., Thompson sampling [40], information directed sampling [41]) do not satisfy these properties. Nevertheless, we believe that the high-level ideas in our proof (see Section 3.1) could be beneficial for future refined analyses of other algorithms.
@@
  • Instance dependent analysis in the frequentist setting. As described in the footnote in Section 3.1, we believe that our analysis does not improve the worst-case regret upper bound in the frequentist setting. On the other hand, our technique can be applied to the instance-dependent analysis [46] for GP-UCB. We expect that our proof strategy could yield a $\widetilde{O}(\sqrt{T})$ instance-dependent regret for GP-UCB by replacing the sample path conditions 3 and 4 in Lemma 2 with the growth condition (Definition 4 in [46]) of the function. It is an interesting direction for future research.
@@ 5 Conclusion
We provide a refined analysis of GP-UCB in the BO problem. For both SE and Matérn kernels, our results improve upon existing regret guarantees and fill the gap between the existing regret of GP-UCB and the current best upper bound in [43]. The core idea of our analysis is to capture the shrinking behavior of the input sequence by relating it to the worst-case upper bound and the sample path regularity conditions. Although our current analysis is limited to GP-UCB in the Bayesian setting, we believe it lays the foundation for several promising future research directions.
@@ A.1 Proof of Lemma 4
[…]
@@
[…]
@@
When $\bm{x}^{\ast}$ is on the boundary of $\mathcal{X}$.
First, when $\bm{x}^{\ast}$ is on the boundary, condition 3 in Lemma 2 holds under $\mathcal{A}$. Then, from the definition of $\mathcal{T}^{c}(\varepsilon)$, $\varepsilon$, and Lemma 20, we have $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ for any $\bm{x}\in\mathcal{X}^{(i)}$. This implies $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\geq c_{\mathrm{lin}}\|\bm{x}-\bm{x}^{\ast}\|_{2}$ from condition 3 in Lemma 2. Since $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\leq\eta_{i}$ from $\mathcal{T}_{i}\subset\mathcal{T}^{c}(\eta_{i})$, we have $\eta_{i}\geq c_{\mathrm{lin}}\|\bm{x}-\bm{x}^{\ast}\|_{2}\Leftrightarrow\eta_{i}c_{\mathrm{lin}}^{-1}\geq\|\bm{x}-\bm{x}^{\ast}\|_{2}$, which implies $\mathcal{X}^{(i)}\subset\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})$. Therefore, when $\bm{x}^{\ast}$ is on the boundary of $\mathcal{X}$, we have
$\gamma_{(T/2^{i-1})}(\mathcal{X}^{(i)})\leq\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})).$ (37)
When $\bm{x}^{\ast}$ is on the interior of $\mathcal{X}$.
Similarly to the boundary case, we consider the case where $\bm{x}^{\ast}$ is on the interior of $\mathcal{X}$. Then, condition 4 in Lemma 2 holds under $\mathcal{A}$. Furthermore, from the definition of $\mathcal{T}^{c}(\varepsilon)$, $\varepsilon$, and Lemma 20, we have $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ for any $\bm{x}\in\mathcal{X}^{(i)}$. This implies $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\geq c_{\mathrm{quad}}\|\bm{x}-\bm{x}^{\ast}\|_{2}^{2}$ from condition 4 in Lemma 2. Since $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\leq\eta_{i}$ from $\mathcal{T}_{i}\subset\mathcal{T}^{c}(\eta_{i})$, we have $\eta_{i}\geq c_{\mathrm{quad}}\|\bm{x}-\bm{x}^{\ast}\|_{2}^{2}\Leftrightarrow\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}}\geq\|\bm{x}-\bm{x}^{\ast}\|_{2}$, which implies $\mathcal{X}^{(i)}\subset\mathcal{B}_{2}(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast})$. Therefore, when $\bm{x}^{\ast}$ is on the interior of $\mathcal{X}$, we have
@@
Finally, combining Eq. (35) with Eqs. (37) and (38), we have
$\begin{split}R_{T}^{(2)}(\varepsilon)&\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)\\ &~{}~{}~{}+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}.\end{split}$ (39)
@@ A.2 Proof of Lemma 5
The existence of $C_{\mathrm{Mat}}$ is guaranteed by the upper bound of MIG established in Corollary 8666If we rely on the result in [54], we can tighten the logarithmic term from $\ln^{\frac{4\nu+d}{2\nu+d}}t$ to $\ln^{\frac{2\nu}{2\nu+d}}t$; however, due to the technical issue of [54] described in Appendix B, we proceed our proof based on Corollary 8. . Note that $C_{\mathrm{Mat}}$ is the constant that may depend on $d,\ell,\nu,r$, and $\sigma^{2}$. Furthermore, we set $\overline{\gamma}_{t}=C_{\mathrm{Mat}}t^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}t$. For function $g(t)\coloneqq\overline{\gamma}_{t}/t$, we have
@@
$\leq\gamma_{t}\left(\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq\eta\}\right)$ (45)
@@
where the second inequality follows from the fact that $k_{\mathrm{Mat\acute{e}rn}}$ is the stationary kernel (namely, $k_{\mathrm{Mat\acute{e}rn}}$ is transition invariant against any shift of inputs). Regarding $\eta_{i}$ in Lemma 4, by setting $T_{i}$ as $T_{i}=T/2^{i-1}$, we have
@@
$\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast}\right)\right)$ (52)
$\leq C_{\mathrm{Mat}}\left(c_{\mathrm{lin}}^{-\frac{2\nu d}{2\nu+d}}\eta_{i}^{\frac{2\nu d}{2\nu+d}}T_{i}^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}T_{i}+\ln^{2}T_{i}\right)$ (53)
$\leq C_{\mathrm{Mat}}\left[c_{\mathrm{lin}}^{-\frac{2\nu d}{2\nu+d}}\widetilde{C}_{\mathrm{Mat}}^{\frac{2\nu d}{2\nu+d}}\beta_{T}^{\frac{\nu d}{2\nu+d}}\left(T_{i}^{-\frac{\nu}{2\nu+d}}\ln^{\frac{4\nu+d}{4\nu+2d}}T_{i}\right)^{\frac{2\nu d}{2\nu+d}}T_{i}^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}T_{i}+\ln^{2}T\right].$ (54)
We further consider the upper bound for the term depending on $T_{i}$: $\left(T_{i}^{-\frac{\nu}{2\nu+d}}\ln^{\frac{4\nu+d}{4\nu+2d}}T_{i}\right)^{\frac{2\nu d}{2\nu+d}}T_{i}^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}T_{i}$. By aligning this term, we have
$\left(T_{i}^{-\frac{\nu}{2\nu+d}}\ln^{\frac{4\nu+d}{4\nu+2d}}T_{i}\right)^{\frac{2\nu d}{2\nu+d}}T_{i}^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}T_{i}$ $=\widetilde{O}\left(T_{i}^{-\frac{2\nu^{2}d}{(2\nu+d)^{2}}+\frac{d}{2\nu+d}}\right)$ (55)
$=\widetilde{O}\left(T_{i}^{\frac{d(2\nu+d)-2\nu^{2}d}{(2\nu+d)^{2}}}\right)$ (56)
$=\widetilde{O}\left(T_{i}^{\frac{d(2\nu+d-2\nu^{2})}{(2\nu+d)^{2}}}\right)$ (57)
$=o(1),$ (58)
where the last line follows from $2\nu+d<2\nu^{2}$, which is implied by the theorem’s condition $2\nu+d\leq\nu^{2}$. Combining the above inequality with Eq. (54), we obtain $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)$. Similarly to the upper bound of $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast}\right)\right)$, we have
@@
From the above inequalities, we have $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)$. By noting $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)$ and $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)$, we have the following inequality for any $i$.
$\sqrt{\max\left\{\gamma_{(T/2^{i-1})}(\mathcal{B}_{2}(\eta_{i}c_{\mathrm{lin}}^{-1};\bm{x}^{\ast})),\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)\right\}}=\widetilde{O}(1).$ (66)
The above inequality and Lemma 4 implies
$R_{T}^{(2)}(\varepsilon)$ $\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+2\overline{i}\sqrt{C\beta_{T}T}\times\widetilde{O}(1)$ (67)
@@
The proof for $k=k_{\mathrm{SE}}$ is not straightforward as with the proof for $k=k_{\mathrm{Mat\acute{e}rn}}$. Specifically, we have to choose a proper $\overline{T}$ so as to obtain an $O(\ln T)$ upper bound of MIG. Let $C_{\mathrm{SE}}>0$ be the constant such that the following inequalities hold:
@@
where $\widetilde{C}_{\mathrm{SE}}>0$ is a sufficiently large constant such that $\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)\geq 4\sqrt{CC_{\mathrm{SE}}\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)+\frac{\pi^{2}}{3T_{i}}$ for any $T_{i}\geq 2$. Hereafter, we define $\overline{\eta}_{i}\coloneqq\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)$. Then, to apply Eq. (70), we consider the lower bound of $T_{i}$ such that $c_{\mathrm{lin}}^{-1}\overline{\eta}_{i}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$ and $\sqrt{c_{\mathrm{quad}}^{-1}\overline{\eta}_{i}}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$ hold. Regarding the condition $c_{\mathrm{lin}}^{-1}\overline{\eta}_{i}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$, we have
$c_{\mathrm{lin}}^{-1}\overline{\eta}_{i}<\sqrt{\frac{2\ell^{2}}{e^{2}c_{d}}}$ $\Leftrightarrow c_{\mathrm{lin}}^{-1}\sqrt{\frac{e^{2}c_{d}}{2\ell^{2}}}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T_{i}<T_{i}^{\frac{1}{2}}$ (76)
$\Leftarrow c_{\mathrm{lin}}^{-1}\sqrt{\frac{e^{2}c_{d}}{2\ell^{2}}}\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\ln^{\frac{d+1}{2}}T<T_{i}^{\frac{1}{2}}$ (77)
$\Leftrightarrow\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}^{2}}{2\ell^{2}c_{\mathrm{lin}}^{2}}\beta_{T}\ln^{d+1}T<T_{i}.$ (78)
Similarly, regarding the condition $\sqrt{c_{\mathrm{quad}}^{-1}\overline{\eta}_{i}}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$, we have
@@
From the above inequalities, we set $\overline{T}$ such that
$\max\left\{\left(\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}}{2\ell^{2}c_{\mathrm{quad}}}\right)^{2},\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}^{2}}{2\ell^{2}c_{\mathrm{lin}}^{2}}\right\}\beta_{T}\ln^{d+1}T<\overline{T}.$ (82)
Then, from $T_{i}\geq\overline{T}$ and Eqs. (70), (76), and (82), we have
$\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(c_{\mathrm{lin}}^{-1}\eta_{i};\bm{x}^{\ast}\right)\right)$ $\leq\gamma_{T_{i}}\left(\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq c_{\mathrm{lin}}^{-1}\eta_{i}\}\right)$ (83)
$\leq C_{\mathrm{SE}}\left(\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{\overline{\eta}_{i}^{2}ec_{d}}\right)}+\ln T_{i}\right)$ (84)
$\leq C_{\mathrm{SE}}\left(\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{\overline{\eta}_{i}^{2}ec_{d}}\right)}+\ln T\right).$ (85)
Furthermore, from $T_{i}\geq\overline{T}$ and Eqs. (70), (79), and (82),
@@
Based on Eqs. (85) and (87), we further consider the lower bound of $T_{i}$ such that
$\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{\overline{\eta}_{i}^{2}ec_{d}}\right)}=O(\ln T)~{}~{}\mathrm{and}~{}~{}\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{quad}}\ell^{2}}{\overline{\eta}_{i}ec_{d}}\right)}=O(\ln T).$ (88)
Regarding the first case in Eq. (88), we have
$\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{\overline{\eta}_{i}^{2}ec_{d}}\geq\sqrt{T_{i}}$ $\Leftrightarrow\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}T_{i}^{-1}\ln^{d+1}T_{i}}\geq\sqrt{T_{i}}$ (89)
$\Leftrightarrow T_{i}^{1/2}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}\ln^{d+1}T_{i}}{2c_{\mathrm{lin}}^{2}\ell^{2}}$ (90)
$\Leftarrow T_{i}^{1/2}\geq\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}\ln^{d+1}T}{2c_{\mathrm{lin}}^{2}\ell^{2}}$ (91)
$\Leftrightarrow T_{i}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}\ln^{d+1}T}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}.$ (92)
Therefore, if $\overline{T}\geq\left(\frac{ec_{d}\widetilde{C}_{\mathrm{SE}}^{2}\beta_{T}\ln^{d+1}T}{2c_{\mathrm{lin}}^{2}\ell^{2}}\right)^{2}$, we have
$C_{\mathrm{SE}}\left(\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{lin}}^{2}\ell^{2}}{\overline{\eta}_{i}^{2}ec_{d}}\right)}+\ln T\right)$ $\leq C_{\mathrm{SE}}\left(\frac{\ln^{d+1}T_{i}}{2^{-d}\ln^{d}T_{i}}+\ln T\right)$ (93)
$\leq C_{\mathrm{SE}}\left(2^{d}\ln T+\ln T\right).$ (94)
Regarding the second case in Eq. (88), we have

### v3

@@ 4 Discussions
  • Optimality. Based on the $\Omega(\sqrt{T})$ lower bound on the expected regret provided by scarlett2018tight, we conjecture that our $\widetilde{O}(\sqrt{T})$ high-probability regret bound for GP-UCB is near-optimal. However, it is not straightforward to extend the lower bound for the expected regret in [scarlett2018tight] to a high probability result. Specifically, the lower bound in [scarlett2018tight] is quantified by a mutual information term (Lemma 4 in [scarlett2018tight]); however, to our knowledge, the technique used to handle this term appears to be specific to the expected regret setting. We believe that the rigorous optimality argument for the Bayesian high probability regret is an important direction for future research.
@@
  • Smoothness condition. In our result for the Matérn kernel, we require an additional smoothness constraint to obtain a $\widetilde{O}(\sqrt{T})$ regret bound [^5] For simplicity, in Theorem 3, we focus on the setting where the resulting regret becomes $\widetilde{O}(\sqrt{T})$ under $k_{\mathrm{Mat\acute{e}rn}}$. This arises the requirement of the additional smoothness condition $2\nu+d\leq\nu^{2}$. On the other hand, we can also apply the same technique even under $2\nu+d>\nu^{2}$. In this case, resulting regret becomes strictly larger than $\widetilde{O}(\sqrt{T})$, while it is strictly smaller than $\widetilde{O}(T^{\frac{\nu+d}{2\nu+d}})$ of the original GP-UCB’s analysis. To overcome this issue in our proof, we believe that we need stronger regularity conditions on the sample path around the maximizer than those assumed in Lemma 2.
@@
  • Extension to the expected regret. Our regret bounds involve regularity constants that depend on the sample path. However, to our knowledge, there is no existing research that rigorously analyzes how these constants depend on the confidence level $\delta_{\mathrm{GP}}$. This makes it difficult to obtain the expected regret guarantees as with the original GP-UCB, whose expected regret bounds are established by properly decreasing the confidence level as a function of $T$ (e.g., [paria2020flexible, takeno2023-randomized]). To overcome this issue, further analysis for Lemma 2, or another idea to quantify the sample path regularities, is required.
@@
  • Extension to other algorithms. One limitation of our technique is its restricted applicability to other algorithms. To apply our proof, at least the algorithm should satisfy the following two conditions: (i) on any index subset, the sub-linear cumulative regret is obtained with high probability (Lemma 21), and (ii) the high probability lenient regret bound is provided (Lemma 6). The existing analysis of the other major algorithms in the Bayesian setting (e.g., Thompson sampling [Russo2014-learning], information directed sampling [russo2014learning]) does not provide these properties. Nevertheless, we believe that the high-level ideas in our proof (see Section 3.1) could be beneficial for future refined analyses of other algorithms.
@@
  • Instance dependent analysis in the frequentist setting. As described in the footnote in Section 3.1, we believe that our analysis does not improve the worst-case regret upper bound in the frequentist setting. On the other hand, our technique can be applied to the instance-dependent analysis [shekhar2022instance] for GP-UCB. We expect that our proof strategy could yield a $\widetilde{O}(\sqrt{T})$ instance-dependent regret for GP-UCB by replacing the sample path condition 3 in Lemma 2 with the growth condition (Definition 4 in [shekhar2022instance]) of the function. It is an interesting direction for future research.
@@ 5 Conclusion
We provide a refined analysis of GP-UCB in the BO problem. For both SE and Matérn kernels, our results improve upon existing regret guarantees and fill the gap between the existing regret of GP-UCB and the current best upper bound in [scarlett2018tight]. The core idea of our analysis is to capture the shrinking behavior of the input sequence by relating it to the worst-case upper bound and the sample path regularity conditions. Although our current analysis is limited to GP-UCB in the Bayesian setting, we believe it lays the foundation for several promising future research directions.
Acknowledgments
We thank Jonathan Scarlett and Shion Takeno for their valuable comments on revising the manuscript.
@@ A.1 Proof of Lemma 4
From Lemma 21, we have the following upper bound for any index set $\mathcal{T}\subset[T]$ under $\mathcal{A}$:
@@
As described in the proof below, these $(\eta_{i})$ are designed so that we can obtain the upper bound of $|\mathcal{T}(\eta_{i})|$ in a dyadic manner. Here, we consider the upper bound of $|\mathcal{T}(\eta_{i})|$ based on the worst-case upper bound in Eq. (15). From the definition of $\mathcal{T}(\eta)$ and Eq. (15) with $\mathcal{T}=[T]$, the condition $|\mathcal{T}(\eta_{1})|\eta_{1}\leq 2\sqrt{C\beta_{T}T\overline{\gamma}_{T}}+\pi^{2}/6$ must be satisfied; otherwise, we have $\sum_{t\in[T]}f(\bm{x}^{\ast})-f(\bm{x}_{t})\geq\sum_{t\in\mathcal{T}(\eta_{1})}f(\bm{x}^{\ast})-f(\bm{x}_{t})\geq|\mathcal{T}(\eta_{1})|\eta_{1}>2\sqrt{C\beta_{T}T\overline{\gamma}_{T}}+\pi^{2}/6$, which contradicts worst-case upper bound in Eq. (15). Therefore, we can obtain the following upper bound:
@@
Conversion of the sub-optimality gap into the upper bound input radius.
From condition 3 in Lemma 2, the definition of $\mathcal{T}^{c}(\varepsilon)$, $\varepsilon$, and Lemma 20, we have $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ for any $\bm{x}\in\mathcal{X}^{(i)}$. This implies $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\geq c_{\mathrm{quad}}\|\bm{x}-\bm{x}^{\ast}\|_{2}^{2}$ from condition 3 in Lemma 2. Since $\forall\bm{x}\in\mathcal{X}^{(i)},f(\bm{x}^{\ast})-f(\bm{x})\leq\eta_{i}$ from $\mathcal{T}_{i}\subset\mathcal{T}^{c}(\eta_{i})$, we have $\eta_{i}\geq c_{\mathrm{quad}}\|\bm{x}-\bm{x}^{\ast}\|_{2}^{2}\Leftrightarrow\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}}\geq\|\bm{x}-\bm{x}^{\ast}\|_{2}$, which implies $\mathcal{X}^{(i)}\subset\mathcal{B}_{2}(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast})$. Therefore, we have
@@
Finally, combining Eq. (35) with Eq. (37), we have
$R_{T}^{(2)}(\varepsilon)\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\max_{i\in[\overline{i}]}\sqrt{\gamma_{(T/2^{i-1})}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)}.$ (38)
@@ A.2 Proof of Lemma 5
The existence of $C_{\mathrm{Mat}}$ is guaranteed by the upper bound of MIG established in Corollary 8666If we rely on the result in [vakili2021information], we can tighten the logarithmic term from $\ln^{\frac{4\nu+d}{2\nu+d}}t$ to $\ln^{\frac{2\nu}{2\nu+d}}t$; however, due to the technical issue of [vakili2021information] described in Appendix B, we proceed our proof based on Corollary 8. . Note that $C_{\mathrm{Mat}}$ is the constant that may depend on $d,\ell,\nu,r$, and $\sigma^{2}$. Furthermore, we set $\overline{\gamma}_{t}=C_{\mathrm{Mat}}t^{\frac{d}{2\nu+d}}\ln^{\frac{4\nu+d}{2\nu+d}}t$. For function $g(t)\coloneqq\overline{\gamma}_{t}/t$, we have
@@
$=\gamma_{t}\left(\{\bm{x}\in\mathbb{R}^{d}\mid\|\bm{x}\|_{2}\leq\eta\}\right)$ (44)
@@
where the second line follows from the fact that $k_{\mathrm{Mat\acute{e}rn}}$ is the stationary kernel (namely, $k_{\mathrm{Mat\acute{e}rn}}$ is transition invariant against any shift of inputs). Regarding $\eta_{i}$ in Lemma 4, by setting $T_{i}$ as $T_{i}=T/2^{i-1}$, we have
@@
From the above inequalities, we have $\gamma_{T/2^{i-1}}\left(\mathcal{B}_{2}\left(\sqrt{c_{\mathrm{quad}}^{-1}\eta_{i}};\bm{x}^{\ast}\right)\right)=\widetilde{O}(1)$. Therefore, Lemma 4 implies
$R_{T}^{(2)}(\varepsilon)$ $\leq 2c_{\mathrm{sup}}\overline{T}+\frac{\pi^{2}}{3}\left(\log_{2}T+1\right)+\frac{2\sqrt{2C\beta_{T}T}}{\sqrt{2}-1}\times\widetilde{O}(1)$ (58)
@@
The proof for $k=k_{\mathrm{SE}}$ is not as straightforward as the proof for $k=k_{\mathrm{Mat\acute{e}rn}}$. Specifically, we have to choose a proper $\overline{T}$ so as to obtain an $O(\ln T)$ upper bound of MIG. Let $C_{\mathrm{SE}}>0$ be the constant such that the following inequalities hold:
@@
where $\widetilde{C}_{\mathrm{SE}}>0$ is a sufficiently large constant such that $\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)\geq 4\sqrt{CC_{\mathrm{SE}}\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)+\frac{\pi^{2}}{3T_{i}}$ for any $T_{i}\geq 2$. Hereafter, we define $\overline{\eta}_{i}\coloneqq\widetilde{C}_{\mathrm{SE}}\sqrt{\beta_{T}}\left(T_{i}^{-\frac{1}{2}}\ln^{\frac{d+1}{2}}T_{i}\right)$. Then, to apply Eq. (61), we consider the lower bound of $T_{i}$ such that $\sqrt{c_{\mathrm{quad}}^{-1}\overline{\eta}_{i}}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$ hold. From the condition $\sqrt{c_{\mathrm{quad}}^{-1}\overline{\eta}_{i}}<\sqrt{2\ell^{2}/(e^{2}c_{d})}$, we have
@@
From the above inequality, we set $\overline{T}$ such that
$\left(\frac{e^{2}c_{d}\widetilde{C}_{\mathrm{SE}}}{2\ell^{2}c_{\mathrm{quad}}}\right)^{2}\beta_{T}\ln^{d+1}T<\overline{T}.$ (70)
Then, from $T_{i}\geq\overline{T}$ and Eqs. (67), and (70),
@@
Based on Eq. (72), we further consider the lower bound of $T_{i}$ such that
$\frac{\ln^{d+1}T_{i}}{\ln^{d}\left(\frac{2c_{\mathrm{quad}}\ell^{2}}{\overline{\eta}_{i}ec_{d}}\right)}=O(\ln T).$ (73)
For the condition in Eq. (73), we have